OpenStudy (anonymous):

Solve for the indicated variable. C=5(F-32)/9 ; for F

OpenStudy (anonymous):

C=5(F-32) / 9 cross multiply first 5(F-32) = 9C distribute next 5F - 160 = 9C add + 160 to both sides 5F = 9C + 160 divide both sides by 5 F = (9/ 5) C + 32

OpenStudy (anonymous):

F = 299792490